How many gallons of a 20% antifreeze solution and a 15% antifreeze solution must be mixed to obtain 25 gal of 17% antifreeze solution? Round to the nearest gallon, if necessary.
____gal of the 20% solution is needed.

Answer by ankor@dixie-net.com(22740)

How many gallons of a 20% antifreeze solution and a 15% antifreeze solution must be mixed to obtain 25 gal of 17% antifreeze solution?
Round to the nearest gallon, if necessary.
:
Let x = amt of 20% solution
the total result is 25 gal, therefore:
(25-x) = amt of 15% solution
:
A typical mixture equation
.20x + .15(25-x) = .17(25)
.20x + 3.75 - .15x = 4.25
.20x - .15x = 4.25 - 3.75
.05x = .50
x = .5/.05
x = 10 gal of the 20% solution
then
25 - 10 = 15 gal of the 15% solution
:
:
Check: .2(10) + .15(15) = .17(25)
2 + 2.25 = 4.25